1. In a study of fatal car accidents, some accidents were attributed to high speed or drunk driving.

a. The probability that the fatality was attributed to high speeds was 0.57.

b. The probability that the fatality was attributed to drunk driving was 0.62.

c. The probability that high speed and drunk driving caused the accident was 0.41.

a. What is the probability that a fatal accident will be attributed to high speeds or drunk driving?

b. What is the probability that a fatal accident will be attributed to neither cause?

c. What is the probability that the fatality will be attributed to drunk driving, but not high speeds?
Answer #1

Since

P(A)=0.57 and P(B)=0.62 also

P(A and B)=0.41

Hence the Venn diagaram is as belowA A B

So, a) P(A or B)= P(A)+ P(B)--(A and B)

Also from Venn diagram

=(0.62+0.57)-0.41= 0.7 8

b) The Probability that fatal accident attributed to neither cause

1-P(A or B)

=1-0.78=

=0.22

C) probability that the fatality will be attributed to drunk not to high speed can be concluded from Venn diagram as

0.62-0.41= 0.21